OpenStudy (anonymous):

Y=5x+4 Determine the slope and y-intercept of the line

OpenStudy (bradely):

y = 5x+4 is in the form y = mx+c m is slope and c is y intercept Source: http://www.mathskey.com/question2answer/

OpenStudy (texaschic101):

In y = mx + b form, the m is the slope and the b is the y intercept EXAMPLE : y = 2x - 8.....slope(m) = 2 and y int (b) = -8

OpenStudy (bradely):

compare both u got answers